Harris Fabrics computes its plantwide predetermined overhead rate annually on the basis of direct labor-hours. At the beginning of the year, it estimated that 38,000 direct labor-hours would be required for the period’s estimated level of production. The company also estimated $507,000 of fixed manufacturing overhead cost for the coming period and variable manufacturing overhead of $2.00 per direct labor-hour. Harris’s actual manufacturing overhead cost for the year was $658,046 and its actual total direct labor was 38,500 hours.
Answer #1

Calculation of predetermined overhead rate

Predetermined overhead rate = Estimated total manufacturing overhead cost / Estimated total amount of the allocation base

Calculation of Estimated variable manufacturing overhead:

It is given that the variable manufacturing overhead is $2 per direct labor-hour and estimated direct labor-hour is 38000 direct labour-hours

Estimated variable manufacturing overhead = Variable manufacturing overhead * Estimated direct labou-hour

= 2*38000

= 76,000

Calculation of total estimated manufacturing overhead:

It is given that the estimated fixed manufacturing overhead is $507000 and the calculated estimated variable manufacturing overhead is $76,000

Total estimated manufacturing overhead = Estimated fixed manufacturing overhead + Estimated variable manufacturing variable manufacturing overhead

= 507000+76000

=583000

Calculation of predetermined overhead rate:

Predetermined overhead rate = 583000/38000

= 15.34